OpenStudy (anonymous):

WILL MEDAL AND FAN!! NEEDS THE ANSWER Write 16/25 as a percentage.

OpenStudy (nathalyn):

divide 16 / 25 and then multiply by 100. 16/ 25 = .64 (x100%) = 64%

OpenStudy (karatechopper):

Well, percentages come out of 100 let's make a proportion \[\frac{ 16 }{ 25}= \frac{ x }{ 100 }\]

OpenStudy (karatechopper):

Whatever you get as x, can be said as the percentage.

OpenStudy (karatechopper):

Cross multiply and divide.

OpenStudy (anonymous):

16/25 = 64/100 = 64%

OpenStudy (anonymous):

ce qui ne signifie "croix multiplient"

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i'm American. i don't understand you

OpenStudy (decentnabeel):

\[\frac{x}{100}=\frac{16}{25}\] \[\mathrm{Multiply\:by\:LCD=}100\] \[\frac{x}{100}100=\frac{16}{25}100\] x=64 answer is .. 64%

OpenStudy (anonymous):

thanks